# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from a Large-Scale Field Experiment

### v3

@@ A.5.2 Computational results for overall planning
Currently, Gurobi adopts three widely-used methods to solve the LP problem, which are respectively primal simplex, dual simplex, and barrier. For OR-Tools, we choose the method of Practical Large-Scale Linear Programming using Primal-Dual Hybrid Gradient (PDLP) (Applegate et al.,, 2021, 2023, 2024), which specializes in solving large-scale LP problems efficiently. Thus, in this experiment, we incorporate four groups (Primal Simplex, Dual Simplex, Barrier, and PDLP) in total, and compare the performance of PDLP to all other groups. For these experiment groups, under different cores and threads, we evaluate their performance in terms of the solving time. Similar to the previous section, we implement this experiment with the Dell PowerEdge R640 computing node. It is noted that in our case, due to the restricted computing resources, the upper limit of the running time is set to be 48 hours ($172,800$ seconds).
@@ A.9 Empirical Evidence on Revenue Independence between SLR and KR
Test 2: Granger causality test. We conduct a Granger causality test (Granger, 1969) to examine whether the sequence of $\textit{RevenueSLR}{i}$ helps forecast $\textit{RevenueKR}{i}$ at various lags, which would indicate a potential dependency. Since our sampled impressions include timestamps, we are able to order them chronologically and construct the required lag structures. As reported in table˜A7, the results reveal no evidence of a significant causal relationship. Across all tested lag lengths (1, 2, 10, and 20), the p-values remain non-significant, suggesting that SLR revenue does not Granger-cause KR revenue.

### v4

@@ A.5.2 Computational results for overall planning
Currently, Gurobi adopts three widely-used methods to solve the LP problem, which are respectively primal simplex, dual simplex, and barrier. For OR-Tools, we choose the method of Practical Large-Scale Linear Programming using Primal-Dual Hybrid Gradient (PDLP) (applegate2021practical; applegate2023faster; applegate2024infeasibility), which specializes in solving large-scale LP problems efficiently. Thus, in this experiment, we incorporate four groups (Primal Simplex, Dual Simplex, Barrier, and PDLP) in total, and compare the performance of PDLP to all other groups. For these experiment groups, under different cores and threads, we evaluate their performance in terms of the solving time. Similar to the previous section, we implement this experiment with the Dell PowerEdge R640 computing node. It is noted that in our case, due to the restricted computing resources, the upper limit of the running time is set to be 48 hours ($172,800$ seconds).
@@ A.9 Empirical Evidence on Revenue Independence between SLR and KR
Test 2: Granger causality test. We conduct a Granger causality test (granger1969investigating) to examine whether the sequence of $\textit{RevenueSLR}{i}$ helps forecast $\textit{RevenueKR}{i}$ at various lags, which would indicate a potential dependency. Since our sampled impressions include timestamps, we are able to order them chronologically and construct the required lag structures. As reported in table˜A7, the results reveal no evidence of a significant causal relationship. Across all tested lag lengths (1, 2, 10, and 20), the p-values remain non-significant, suggesting that SLR revenue does not Granger-cause KR revenue.